# PS3C: An Ensemble-Based Two-Step Framework for Classification of Pep Smear Cell Images

### v1

@@
PS3C: An Ensemble-Based Two-Step Framework for Classification of Pep Smear Cell Images
Early detection of cervical cancer is crucial for improving patient outcomes and reducing mortality by identifying precancerous lesions as soon as possible. As a result, the use of pap smear screening has significantly increased, leading to a growing demand for automated tools that can assist cytologists managing their rising workload. To address this, the Pep Smear Cell Classification Challenge (PS3C) has been organized in association with ISBI in 2025. This project aims to promote the development of automated tools for pep smear images classification. The analyzed images are grouped into four categories: healthy, unhealthy, both, and rubbish images which are considered as unsuitable for diagnosis. In this work, we propose a two-stage ensemble approach: first, a neural network determines whether an image is rubbish or not. If not, a second neural network classifies the image as containing a healthy cell, an unhealthy cell, or both.
@@ 1 Introduction
Cervical cancer remains one of the most prevalent cancers among women, causing over 300,000 deaths annually [1]. The widespread adoption of Pap smear screening has significantly improved early detection of cancerous lesions [2, 3]. This process involves collecting cervical cell samples, preparing them as smears on glass slides, and digitizing them using 3D Scanners. The resulting high-resolution images are then divided into smaller patches [4], preserving regions likely to contain diagnostically relevant cells, as illustrated in Figure 1. Cytologists analyze these slices to identify unhealthy cells, but the vast number of images makes this task time-consuming, resource-intensive, and highly dependent on practitioner expertise [5]. Deep learning models [6] offer a promising avenue to enhance this process and support cytologists to manage their increasing workload by classifying cells as healthy or unhealthy. In medical imaging, extensive efforts have been dedicated to develop deep learning methods for various tasks in cell analysis, including segmentation [7, 8], detection [9, 10], and classification [11, 12]. However, Pap smear cell classification remains challenging due to the limited number of publicy available dataset [13, 14, 15], the presence of images unsuitable for evaluation (e.g., artifacts, poor resolution) and the class imbalance, where unhealthy cells are significantly outnumbered by healthy ones, as illustrated by Figure 2. To address this, the PS3C Challenge introduced the APACC dataset [5] to facilitate the development and evaluation of algorithms capable of classifying pep smell images.
@@ 3.2 Step 1: Binary classification task
The first step of our framework involves predicting whether an image should be classified as rubbish or not, formulated as a binary classification task. During training, we consider two labels: rubbish and non-rubbish (where healthy, unhealthy, and both are grouped under the non-rubbish label). Each image $x\in\mathbb{R}^{3\times N\times N}$ is processed by a backbone network, denoted as $\Phi_{1}:\mathbb{R}^{3\times N\times N}\to\mathbb{R}^{d}$, pre-trained on ImageNet [noauthor_imagenet_nodate], to extract a feature representation $h\in\mathbb{R}^{d}$, such that:
@@ 4 Experimental setup
For the first and second steps, each model was trained with a batch size of $32$, using the AdamW [loshchilov_decoupled_2019] optimizer for 80 epochs, with a learning rate of $10^{-5}$. The training required a GPU with 80GB of memory.

### v2

# An Ensemble-Based Two-Step Framework for Classification of Pap Smear Cell Images

@@
An Ensemble-Based Two-Step Framework for Classification of Pap Smear Cell Images
Early detection of cervical cancer is crucial for improving patient outcomes and reducing mortality by identifying precancerous lesions as soon as possible. As a result, the use of pap smear screening has significantly increased, leading to a growing demand for automated tools that can assist cytologists managing their rising workload. To address this, the Pap Smear Cell Classification Challenge (PS3C) has been organized in association with ISBI in 2025. This project aims to promote the development of automated tools for pap smear images classification. The analyzed images are grouped into four categories: healthy, unhealthy, both, and rubbish images which are considered as unsuitable for diagnosis. In this work, we propose a two-stage ensemble approach: first, a neural network determines whether an image is rubbish or not. If not, a second neural network classifies the image as containing a healthy cell, an unhealthy cell, or both.
@@ 1 Introduction
Cervical cancer remains one of the most prevalent cancers among women, causing over 300,000 deaths annually [1]. The widespread adoption of Pap smear screening has significantly improved early detection of cancerous lesions [2, 3]. This process involves collecting cervical cell samples, preparing them as smears on glass slides, and digitizing them using 3D Scanners. The resulting high-resolution images are then divided into smaller patches [4], preserving regions likely to contain diagnostically relevant cells, as illustrated in Figure 1. Cytologists analyze these slices to identify unhealthy cells, but the vast number of images makes this task time-consuming, resource-intensive, and highly dependent on practitioner expertise [5]. Deep learning models [6] offer a promising avenue to enhance this process and support cytologists to manage their increasing workload by classifying cells as healthy or unhealthy. In medical imaging, extensive efforts have been dedicated to develop deep learning methods for various tasks in cell analysis, including segmentation [7, 8], detection [9, 10], and classification [11, 12]. However, Pap smear cell classification remains challenging due to the limited number of publicy available dataset [13, 14, 15], the presence of images unsuitable for evaluation (e.g., artifacts, poor resolution) and the class imbalance, where unhealthy cells are significantly outnumbered by healthy ones, as illustrated by Figure 2. To address this, the PS3C Challenge introduced the APACC dataset [5] to facilitate the development and evaluation of algorithms capable of classifying pap smell images.
@@ 3.2 Step 1: Binary classification task
The first step of our framework involves predicting whether an image should be classified as rubbish or not, formulated as a binary classification task. During training, we consider two labels: rubbish and non-rubbish (where healthy, unhealthy, and both are grouped under the non-rubbish label). Each image $x\in\mathbb{R}^{3\times N\times N}$ is processed by a backbone network, denoted as $\Phi_{1}:\mathbb{R}^{3\times N\times N}\to\mathbb{R}^{d}$, pre-trained on ImageNet [46], to extract a feature representation $h\in\mathbb{R}^{d}$, such that:
@@ 4 Experimental setup
For the first and second steps, each model was trained with a batch size of $32$, using the AdamW [47] optimizer for 80 epochs, with a learning rate of $10^{-5}$. The training required a GPU with 80GB of memory.